# Fano varieties of middle pseudoindex

### v1

@@ 1. Introduction
[…]
@@ 2.1. Fano varieties with large pseudoindex
Let $X$ be a smooth Fano variety of dimension $n$ with pseudoindex $\iota_{X}$. Then the following holds.
@@
By [2, Lemme 2.5 (a)], $W$ is a smooth Fano variety whose pseudoindex is at least $\iota_{X}=\dfrac{n+1}{2}$. Applying Theorem 2.1, $W$ is isomorphic to ${\mathbb{P}}^{\frac{n-1}{2}}$, ${\mathbb{P}}^{\frac{n+1}{2}}$ or $Q^{\frac{n+1}{2}}$. By [5, Proposition 4.3], there exists a vector bundle ${\mathcal{E}}$ over $W$ such that $X\cong{\mathbb{P}}({\mathcal{E}})$. When $\pi:X\to W$ is a ${\mathbb{P}}^{\frac{n+1}{2}}$-bundle, $W$ is isomorphic to ${\mathbb{P}}^{\frac{n-1}{2}}$. In this case, [2, Lemme 2.5 (c)] and [14] tell us that $X$ is isomorphic to ${\mathbb{P}}^{\iota_{X}-1}\times{\mathbb{P}}^{\iota_{X}}$. When $\pi:X\to W$ is a ${\mathbb{P}}^{\frac{n-1}{2}}$-bundle, our assertion is derived from [5, Corollary 4.7]. Thus our assertion holds.
@@ 2.2. Extremal contractions
Let $X$ be a smooth projective variety, and let $\varphi:X\to Y$ be a contraction of an extremal ray $R$ such that any fiber has dimension $d$ and $\ell(R)=d+1$. Then $\varphi$ is a projective bundle.
@@ 2.3. Families of Rational Curves
Moreover if ${\mathcal{M}}$ is unsplit, this inequality holds for any point $x\in{\rm Locus}({\mathcal{M}})$.
@@ 3.1. The case when $X$ admits a birational elementary contraction
Let us consider a minimal rational component $\mathcal{M}$ on $X$. According to Theorem 2.7, the anticanonical degree of $\mathcal{M}$ is at most $n$. Combining with our assumption that $\iota_{X}=\frac{n+1}{2}$ and [10, II, Proposition 2.2], it follows that $\mathcal{M}$ is an unsplit covering family. For any $x\in F$, Proposition 2.8 implies $\dim{\rm Locus}(\mathcal{M}_{x})\geq\frac{n-1}{2}$. To establish our assertion, let us assume the contrary, namely $\text{codim}_{X}E\geq 2$. Then, by Theorem 2.4, it follows that $\dim F\geq\frac{n+3}{2}$. Consequently, $\dim({\rm Locus}(\mathcal{M}_{x})\cap F)\geq 1$. This leads to a contradiction.
@@
To prove this proposition, throughout this subsection, we stay within the confines of the present subsection, maintaining the setting where $X$ is a smooth Fano variety with $\iota_{X}=\dfrac{n+1}{2}$ and $\rho_{X}>1$. Assume that any elementary contraction of $X$ is of fiber type. For different extremal rays $R$ and $R^{\prime}$ of $\overline{NE}(X)$, consider the elementary contractions $\varphi:X\to Y$ and $\psi:X\to Z$ associated to $R$ and $R^{\prime}$ respectively. We denote by $F_{\rm gen}$ (resp. $F_{\rm gen}^{\prime}$) any fiber of $\varphi$ (resp. $\psi$) whose dimension is equal to $\dim X-\dim Y$ (resp. $\dim X-\dim Z$). Using Theorem 2.4, we infer
(3) $\dim F_{\rm gen}\geq\ell(R)-1\geq\dfrac{n-1}{2}\quad\mbox{and}\quad\dim F_{\rm gen}^{\prime}\geq\ell(R^{\prime})-1\geq\dfrac{n-1}{2}.$
@@
(4) $\dim F_{\rm gen}+\dim F_{\rm gen}^{\prime}-n\leq 0,$
it turns out that $\dim F_{\rm gen}$ and $\dim F_{\rm gen}^{\prime}$ are at most $\dfrac{n+1}{2}$. Thus $(\dim F_{\rm gen},\dim Y)$ and $(\dim F_{\rm gen}^{\prime},\dim Z)$ are either:

### v2

@@ 1. Introduction
    ${\mathbb{P}}({\mathcal{O}}_{{\mathbb{P}}^{i_{X}}}(2)\oplus{\mathcal{O}}_{{\mathbb{P}}^{i_{X}}}(1)^{\oplus i_{X}-1}),\quad{\mathbb{P}}^{i_{X}-1}\times Q^{i_{X}}\quad\mbox{or}\quad{\mathbb{P}}(T_{{\mathbb{P}}^{i_{X}}}).$
@@ 2.1. Fano varieties with large pseudoindex
Let $X$ be a smooth Fano variety of dimension $n$ with pseudoindex $\iota_{X}$. Then, the following holds.
@@
By [2, Lemme 2.5 (a)], $W$ is a smooth Fano variety whose pseudoindex is at least $\iota_{X}=\dfrac{n+1}{2}$. Applying Theorem 2.1, $W$ is isomorphic to ${\mathbb{P}}^{\frac{n-1}{2}}$, ${\mathbb{P}}^{\frac{n+1}{2}}$ or $Q^{\frac{n+1}{2}}$. By [5, Proposition 4.3], there exists a vector bundle ${\mathcal{E}}$ over $W$ such that $X\cong{\mathbb{P}}({\mathcal{E}})$. When $\pi:X\to W$ is a ${\mathbb{P}}^{\frac{n+1}{2}}$-bundle, $W$ is isomorphic to ${\mathbb{P}}^{\frac{n-1}{2}}$. In this case, [2, Lemme 2.5 (c)] and [14] tell us that $X$ is isomorphic to ${\mathbb{P}}^{\iota_{X}-1}\times{\mathbb{P}}^{\iota_{X}}$. When $\pi:X\to W$ is a ${\mathbb{P}}^{\frac{n-1}{2}}$-bundle, our assertion is derived from [5, Corollary 4.7]. Thus, our assertion holds.
@@ 2.2. Extremal contractions
Let $X$ be a smooth projective variety, and let $\varphi:X\to Y$ be a contraction of an extremal ray $R$ such that any fiber has dimension $d$ and $\ell(R)=d+1$. Then, $\varphi$ is a projective bundle.
@@
Remark 2.7.
For a smooth projective variety $X$, let $\varphi:X\to Y$ and $\psi:X\to Z$ be different elementary contractions of $X$. Then the fibers of $\varphi$ and $\psi$ have a finite intersection. We use this property several times in this paper.
@@ 2.3. Families of Rational Curves
Moreover, if ${\mathcal{M}}$ is unsplit, this inequality holds for any point $x\in{\rm Locus}({\mathcal{M}})$.
@@ 3.1. The case when $X$ admits a birational elementary contraction
Let us consider a minimal rational component $\mathcal{M}$ on $X$. According to Theorem 2.8, the anticanonical degree of $\mathcal{M}$ is at most $n$. Combining with our assumption that $\iota_{X}=\frac{n+1}{2}$ and [10, II, Proposition 2.2], it follows that $\mathcal{M}$ is an unsplit covering family. For any $x\in F$, Proposition 2.9 implies $\dim{\rm Locus}(\mathcal{M}_{x})\geq\frac{n-1}{2}$. To establish our assertion, let us assume the contrary, namely $\text{codim}_{X}E\geq 2$. Then, by Theorem 2.4, it follows that $\dim F\geq\frac{n+3}{2}$. Consequently, $\dim({\rm Locus}(\mathcal{M}_{x})\cap F)\geq 1$. By [10, II, Corollary 4.21], this leads to a contradiction.
@@
To prove this proposition, throughout this subsection, we stay within the confines of the present subsection, maintaining the setting where $X$ is a smooth Fano variety with $\iota_{X}=\dfrac{n+1}{2}$ and $\rho_{X}>1$. Assume that any elementary contraction of $X$ is of fiber type. For different extremal rays $R$ and $R^{\prime}$ of $\overline{NE}(X)$, consider the elementary contractions $\varphi:X\to Y$ and $\psi:X\to Z$ associated to $R$ and $R^{\prime}$ respectively. We denote by $F$ (resp. $F^{\prime}$) any fiber of $\varphi$ (resp. $\psi$). Using Theorem 2.4, we infer
(3) $\dim F\geq\ell(R)-1\geq\dfrac{n-1}{2}\quad\mbox{and}\quad\dim F^{\prime}\geq\ell(R^{\prime})-1\geq\dfrac{n-1}{2}.$
@@
(4) $\dim F+\dim F^{\prime}-n\leq 0,$
it turns out that $\dim F$ and $\dim F^{\prime}$ are at most $\dfrac{n+1}{2}$. Thus, denoting by $F_{\rm gen}$ (resp. $F_{\rm gen}^{\prime}$) any fiber of $\varphi$ (resp. $\psi$) whose dimension is equal to $\dim X-\dim Y$ (resp. $\dim X-\dim Z$), $(\dim F_{\rm gen},\dim Y)$ and $(\dim F_{\rm gen}^{\prime},\dim Z)$ are either:
@@ Acknowledgments
Conflict of Interest.
The author has no conflicts of interest directly relevant to the content of this article.
Data availability.
Data sharing not applicable to this article as no data sets were generated or analyzed during the current study.